OpenStudy (anonymous):

You decide to put $100 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds?

OpenStudy (anonymous):

Please help!! Will give a medal!!!

OpenStudy (rainbow_rocks03):

plz help

OpenStudy (anonymous):

Do you know it? @rainbow_rocks03

OpenStudy (rainbow_rocks03):

no sorry I don't

OpenStudy (anonymous):

Do the interest on on one whole year and then just multiply until you get 3000 or more and you'll find your answer.

OpenStudy (anonymous):

@PrincessRu How do you do that? These are my answer choices. I think the third one is right. 170.202 years 14.3129 years 171.755 years 168.354 years

OpenStudy (anonymous):

Wait now I am getting A

OpenStudy (anonymous):

But I am not a 100% postive at all.

OpenStudy (tkhunny):

Can you do the first month?

OpenStudy (anonymous):

@tkhunny Would you do 100 times .02?

OpenStudy (tkhunny):

.02 = 2% and is an annual rate. You need a monthly rate.

OpenStudy (anonymous):

So how do I do that? @tkhunny

OpenStudy (tkhunny):

You'll kick yourself. 0.02 / 12 = 0.00166666...

OpenStudy (anonymous):

hahaa! so then do I take that decimal and multipy it by something to get the years? @tkhunny

OpenStudy (tkhunny):

100 * 0.00166666 = 17¢ That's just the interest. You need to accumulate all the interest and the initial deposit. Annual Interest: 0.02 Monthly Interest: 0.02 / 12 Monthly Accumulation: 1 + (0.02 / 12) You task is to see how long this takes to get to $3000 \(100 [1 + (0.02/12)]^{n} = 3000\) Use your best logarithms to solve for 'n'.

OpenStudy (anonymous):

Wow I would have never gotten to that. I am not sure how to solve that. @tkhunny

OpenStudy (anonymous):

@tkhunny i got it to log29(.02/12)=n